Mrs. Garcia invests a total of $7214 into two savings accounts. One account yields 8.5% simple interest and the other 9.5% simple interest. She would like to find the amount placed in each account if a total of $624.13 in interest is received after one year. Write a system of linear equations to express this
Answer by richwmiller(17219)

Total amount of money invested: $7214
x+y=7214,
Total yearly interest for the two accounts is: $624.13
0.085*x+0.095*y=624.13
x=7214-y
Substitute for x
0.085*(7214-y)+0.095*y=624.13
Multiply out
613.19-0.085y+0.095*y=624.13
Combine like terms.
0.01*y=10.94
Isolate y
y=$ 1094.00 at 9.5%
x=7214-y
Calculate x
x=$ 6120.00 at 8.5%
Check
0.085*6120.0+0.095*1094.0=624.13
520.2+103.93=624.13
624.13=624.13
If this statement is TRUE and neither amount is negative then all is well